### v1

# Martingale Approach to Gambler’s Ruin Problem for Correlated Random Walks

Abstract The gambler’s ruin problem for correlated random walks, both with and without delays, is addressed by Optional Stopping Theorem for martingales.
Keywords: Gambler’s Ruin, Martingale, Optional Stopping Theorem, Correlated Random Walk
2020 Mathematics Subject Classification: 60G42, 60J10
1 A Two-Pattern Problem
A casino flips a fair coin. A gambler gets \$1, if he sees HH (including overlapping occurrences, that is, for HHH he will get \$2). He loses a dollar, if it is TH. Here are some facts that we know about this game:
  1. 1.
    The expected number of occurrences of HH and TH in $n$ trials is $(n-1)/4$ for both patterns.
  2. 2.
    The expected waiting time until the first HH is 6.
  3. 3.
    The expected waiting time until the first TH is 4.
  4. 4.
    The expected waiting time until the first occurrence of either pattern is 3.
  5. 5.
    The probability that HH will occur before TH is .25.
The first statement is a trivial fact that one can get by taking the expectation of a sum of appropriate indicators. The other statements are more interesting. There is a long list of articles on this topic, the occurrence of patterns. An approach based on generating functions can be found in Feller (1968). In a private communication, Donald Knuth indicated to us that the key observation—that the expected waiting time until the first occurrence of a pattern depends on self-overlapping of the pattern with itself—was first made in Solov’ev (1966). A combinatorial method was developed in Guibas and Odlyzko (1981). More relevant to our approach is a nice martingale technique introduced in Li (1980).
Now, let us ask a different question. Let $A$, $B$ be positive integers. What is the probability of getting to \$$A$ before reaching -\$$B$? This is a variation of the famous gambler’s ruin problem for random walks that one can find in classical texts such as Feller (1968). However, here we have two important differences from the standard setup: (1) the payments are not independent, and (2) the payments take three values (1, 0, and -1).
2 Correlated Random Walk
Consider a two-state Markov Chain $\{X_{n}\}_{n\geq 1}$ with state set $\{1,-1\}$, transition matrix
$\begin{pmatrix}p&1-p\\ 1-q&q\\ \end{pmatrix},$
and initial distribution $(p_{0},1-p_{0})$, where $10<p,q<$ 1 and $0\leq p_{0}\leq 1$. In what follows, we assume that the initial distribution is always a stationary one, that is, in this case, $p_{0}=(1-q)/(2-p-q)$. This assumption helps to eliminate the extra variable, $p_{0}$, in our formulas. But all our derivations can be easily adjusted to an arbitrary initial distribution.
Let $S_{n}=X_{1}+\cdots+X_{n}$ denote a correlated random walk or CRW, and let ${\mathcal{F}}_{n}=\sigma(X_{1},\dots,X_{n})$, the $\sigma$-field generated by $\{X_{n}\}_{n\geq 1}$. For any integers $A,B>0$, let us consider the following stopping time (w.r.t to the filtration $\{{\mathcal{F}}_{n}\}_{n\geq 1}$):
$\tau=\inf\{k:S_{k}=A\mbox{ or }S_{k}=-B\}.$
One important technical observation is that ${\mathbf{E}}(\tau)<\infty$. It is a well-known fact, so we will skip the proof. Then the “ruin” probabilities are: $\alpha={\mathbf{P}}(S_{\tau}=A)$ and $\beta={\mathbf{P}}(S_{\tau}=-B)$.
When $p=1-q$, $\{X_{n}\}_{n\geq 1}$ are independent identically distributed (iid) random variables. The gambler’s ruin problem can be solved with the help of various techniques. The most popular ones are first-step analysis and the martingale approach. A review of both methods can be found in Steele (2001). In particular, the following martingales are useful. For a symmetric random walk ($p=q=1/2$), they are $S_{n}$ and $S_{n}^{2}-n$. For asymmetric random walks, they are $(q/p)^{S_{n}}$ and $S_{n}-(p-q)n$. Then by a standard application of the Optional Stopping Theorem (OST) one can show that, for example, in the symmetric case $\alpha=B/(A+B)$ and ${\mathbf{E}}(\tau)=AB$.
The general Markovian case was first treated in Mohan (1955) with the help of the first-step analysis. To the best of our knowledge, nobody has used a martingale approach to address the gambler’s ruin problem for correlated random walks. So, our first task is to re-derive the formulas from Mohan (1955) with the help of the Optional Stopping Theorem. For that, we need to construct an additive martingale (similar to $S_{n}-(p-q)n$ in the iid case).
The starting point is a formula for the conditional expectation ${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})$. Observe first that
${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}(X_{n}|X_{n-1}=1)\mathbb{I}_{X_{n-1}=1}+{\mathbf{E}}(X_{n}|X_{n-1}=-1)\mathbb{I}_{X_{n-1}=-1}$
$=$ $\big{[}1\cdot p+(-1)\cdot(1-p)\big{]}\frac{X_{n-1}+1}{2}+\big{[}1\cdot(1-q)+(-1)\cdot q\big{]}\frac{1-X_{n-1}}{2}$
$=$ $(p+q-1)X_{n-1}+(p-q).$
Next, if we define $Y_{1}=X_{1}$ and $Y_{n}=X_{n}-(p+q-1)X_{n-1}-(p-q)$ for $n\geq 2$, then
$M_{n}=$ $Y_{1}+\cdots+Y_{n}$
$=$ $(2-p-q)S_{n-1}+X_{n}-(p-q)(n-1)$ (1)
$=$ $(2-p-q)S_{n}+(p+q-1)X_{n}-(p-q)(n-1)$
is a martingale w.r.t. the filtration $\{{\mathcal{F}}_{n}\}_{n\geq 1}$ as a sum of martingale-differences.
2.1 Ruin Probability for Symmetric CRW
If $p=q$, then
$M_{n}=(2-2p)S_{n}+(2p-1)X_{n}.$
The Optional Stopping Theorem immediately gives us that
$0={\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=(2-2p){\mathbf{E}}(S_{\tau})+(2p-1){\mathbf{E}}(X_{\tau})$
or
$0=(2-2p)(\alpha A-\beta B)+(2p-1)(\alpha-\beta).$
Since we also have $\alpha+\beta=1$, we immediately get
$\alpha=\frac{B-1+\frac{1}{2}\frac{1}{1-p}}{A+B-2+\frac{1}{1-p}}.$ (2)
Note that we have intuitive answers for special cases:
  1. 1.
    if $A=B$ (symmetry), then $\alpha=1/2$,
  2. 2.
    if $p=1/2$ (iid), then $\alpha=B/(A+B)$,
  3. 3.
    if $p\to 1$, then $\alpha\to 1/2$ (everything is decided by the first flip)
2.2 Expected Waiting Time until Ruin for Symmetric CRW
The iid case tells us that we should try to find a “quadratic” martingale. This suggests the following candidate:
$M_{n}=S_{n}^{2}+aS_{n}X_{n}-bn.$
If it is possible to find such values of $a$ and $b$ so $M_{n}$ forms a martingale, then the OST gives us
$(1+a)-b={\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=A^{2}\alpha+B^{2}\beta+a(A\alpha+B\beta)-b{\mathbf{E}}(\tau).$
Since we already know $\alpha$ and $\beta$, we have a formula for the expected waiting time
${\mathbf{E}}(\tau)=\frac{1}{b}\big{[}b-(1+a)+A^{2}\alpha+B^{2}\beta+a(A\alpha+B\beta)\big{]}.$ (3)
Now, using ${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})=(2p-1)X_{n-1}$ and $X_{n}^{2}=1$ we get that
${\mathbf{E}}(M_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}(S_{n}^{2}+aS_{n}X_{n}-bn|{\mathcal{F}}_{n-1})$
$=$ ${\mathbf{E}}((S_{n-1}+X_{n})^{2}+a(S_{n-1}+X_{n})X_{n}-bn|{\mathcal{F}}_{n-1})$
$=$ $S_{n-1}^{2}+\big{[}2(2p-1)+a(2p-1)\big{]}S_{n-1}X_{n-1}+\big{[}1+a-bn\big{]}.$
Therefore, to make it a martingale we need $a$ and $b$ to satisfy the following system of equations:
$\left\{\begin{array}[]{l}2(2p-1)+a(2p-1)=a\\ 1+a-bn=-b(n-1)\end{array}\right.$
This gives us $a=(2p-1)/(1-p)$ and $b=p/(1-p)$. Plugging these values into (3), after some algebra, we obtain that
${\mathbf{E}}(\tau)=AB\frac{1-p}{p}+(A+B)\frac{2p-1}{2p}.$ (4)
Here are some special cases:
  1. 1.
    if $p=1/2$ (iid), then ${\mathbf{E}}(\tau)=AB$,
  2. 2.
    if $p=1$, then ${\mathbf{E}}(\tau)=(A+B)/2$ (going straight to $A$ or $-B$ depending on initial value $X_{1}$),
  3. 3.
    if $A=B$, then ${\mathbf{E}}(\tau)=A^{2}\frac{1-p}{p}+A\frac{2p-1}{p}$.
2.3 Ruin Probability for Asymmetric CRW
The iid case together with our first martingale construction suggests an “exponential” martingale
$M_{n}=\lambda^{S_{n}+aX_{n}},$
where $\lambda>0$ and $a\in\mathbb{R}$. Again, let us assume first that we are successful in finding $\lambda>0$ and $a$ that make $M_{n}$ a martingale. Then in addition to $\alpha+\beta=1$, the OST gives another linear equation:
${\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=\lambda^{A+a}\alpha+\lambda^{-B-a}\beta.$ (5)
So, let us figure out these $\lambda$ and $a$. First, we have that
${\mathbf{E}}(M_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}(\lambda^{S_{n-1}+(1+a)X_{n}+aX_{n-1}-aX_{n-1}}|{\mathcal{F}}_{n-1})=M_{n-1}{\mathbf{E}}(\lambda^{(1+a)X_{n}-aX_{n-1}}|{\mathcal{F}}_{n-1}).$
Can we make the conditional expectation in the last expression to be equal to 1? Yes. As before,
${\mathbf{E}}(\lambda^{(1+a)X_{n}-aX_{n-1}}|{\mathcal{F}}_{n-1})=$ $\mathbin{\hphantom{+}}\big{[}p\lambda^{(1+a)-a}+(1-p)\lambda^{-(1+a)-a}\big{]}\frac{X_{n-1}+1}{2}$
$+\big{[}(1-q)\lambda^{(1+a)+a}+q\lambda^{-(1+a)+a}\big{]}\frac{1-X_{n-1}}{2}$
$=$ $\mathbin{\hphantom{+}}\big{[}p\lambda+(1-p)\lambda^{-1-2a}-(1-q)\lambda^{1+2a}-q\lambda^{-1}\big{]}\frac{X_{n-1}}{2}$
$+\big{[}p\lambda+(1-p)\lambda^{-1-2a}+(1-q)\lambda^{1+2a}+q\lambda^{-1}\big{]}\frac{1}{2}.$
Next, equating the first coefficient to 0 and the second one to 2, we get the following non-linear system of equations for $\lambda$ and $a$
$\left\{\begin{array}[]{l}p\lambda+(1-p)\lambda^{-1-2a}=1\\ (1-q)\lambda^{1+2a}+q\lambda^{-1}=1\end{array}\right.$
Because of the iid martingale $(q/p)^{S_{n}}$, it is natural to try $\lambda=q/p$. And it works, $\lambda=q/p$ and
$a=\frac{1}{2}\left[\log_{q/p}\frac{1-p}{1-q}-1\right]$
is a solution. Note that $a$ itself is not really a part of equation (5), it enters as
$\left(\frac{q}{p}\right)^{a}=\sqrt{\frac{(1-p)p}{(1-q)q}}.$
Since our Markov chain starts from the stationary distribution
${\mathbf{E}}(M_{1})=$ $\left(\frac{q}{p}\right)^{1+a}\frac{1-q}{2-p-q}+\left(\frac{q}{p}\right)^{-1-a}\frac{1-p}{2-p-q}=\frac{\sqrt{(1-q)(1-p)}}{2-p-q}\left[\sqrt{\frac{q}{p}}+\sqrt{\frac{p}{q}}\right]$
Finally, plugging it into (5) and solving with respect $\alpha$ gives us
$\alpha=\frac{\left(\frac{q}{p}\right)^{-B}\sqrt{\frac{(1-q)q}{(1-p)p}}-\frac{\sqrt{(1-q)(1-p)}}{2-p-q}\left[\sqrt{\frac{q}{p}}+\sqrt{\frac{p}{q}}\right]}{\left(\frac{q}{p}\right)^{-B}\sqrt{\frac{(1-q)q}{(1-p)p}}-\left(\frac{q}{p}\right)^{A}\sqrt{\frac{(1-p)p}{(1-q)q}}}.$ (6)
If $p=1-q$, $p\neq 1/2$ , then formula (6) simplifies to a well-known expression for the asymmetric iid case (for example, (Steele, 2001)).
2.4 Expected Waiting Time until Ruin for Asymmetric CRW
Everything is ready now. Applying the OST to the martingale (2), we have the following equation for ${\mathbf{E}}(\tau)$:
${\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=(2-p-q)(A\alpha-B\beta)+(p+q-1)(\alpha-\beta)-(p-q)({\mathbf{E}}(\tau)-1).$
Since we already know $\alpha$ (and $\beta=1-\alpha$) and ${\mathbf{E}}(M_{1})={\mathbf{E}}(X_{1})=(p-q)/(2-p-q)$, the formula for ${\mathbf{E}}(\tau)$ immediately follows.
3 Correlated Random Walk with Delays
As we mentioned in the introduction, accounting for the Markovian dependence is not enough to solve our two-pattern problem. The payments are quite often equal to 0. Random walks associated with such random variables are called “random walks with delays”. The gambler’s ruin problem with delays for the iid case was solved in Gut (2013) with the help of martingales and the OST. Here we extend these results to CRW with delays.
3.1 Ruin Probability for Symmetric CRW with Delays
Consider a three-state Markov Chain $\{X_{n}\}_{n\geq 1}$ with state set $\{1,0,-1\}$ and transition matrix
$\begin{pmatrix}p&r&1-p-r\\ (1-q)/2&q&(1-q)/2\\ 1-p-r&r&p\\ \end{pmatrix},$
where $10\leq p,q,r<$ 1 and $0\leq p+r\leq 1$. We assume that the initial distribution is stationary, that is, it is given by
$\left(\frac{1-q}{2(1+r-q)},\frac{r}{(1+r-q)},\frac{1-q}{2(1+r-q)}\right).$
We will keep the same notation for random walks, ruin probabilities, filtration and stopping time. We call this random walk symmetric, because states 1 and -1 enter into the picture in a symmetric way, they are interchangeable.
As before, the first step is to figure out the conditional expectation ${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})$. We will use a technique similar to the one employed in the previous section. But note that in this case, $X_{n}^{2}$ is still a random variable, and the expressions for indicators of events $\{X_{n}=k\}$, $k=1,0,-1$ now depend on both $X_{n}$ and $X_{n}^{2}$. More specifically, we have
$\mathbb{I}_{X_{n}=1}=(X_{n}^{2}+X_{n})/2,\quad\quad\mathbb{I}_{X_{n}=0}=1-X_{n}^{2},\quad\quad\mathbb{I}_{X_{n}=-1}=(X_{n}^{2}-X_{n})/2.$
As a consequence, we get
${\mathbf{E}}(X_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}(X_{n}|X_{n-1}=1)\mathbb{I}_{X_{n-1}=1}+{\mathbf{E}}(X_{n}|X_{n-1}=0)\mathbb{I}_{X_{n-1}=0}+{\mathbf{E}}(X_{n}|X_{n-1}=-1)\mathbb{I}_{X_{n-1}=-1}$
$=$ $\mathbin{\hphantom{+}}\big{[}p-(1-p-r)\big{]}(X_{n-1}^{2}+X_{n-1})/2$
$+\big{[}(1-q)/2-(1-q)/2\big{]}(1-X_{n-1}^{2})$
$+\big{[}(1-p-r)-p\big{]}(X_{n-1}^{2}-X_{n-1})/2$
$=$ $\mathbin{\hphantom{+}}(2p+r-1)X_{n-1}.$
Note that because of the symmetry of the transition matrix, $X_{n-1}^{2}$ is not a part of the formula.
So, a natural candidate for another “useful” martingale is
$M_{n}=S_{n}+aX_{n}=S_{n-1}+(a+1)X_{n}.$
Using the expression for the conditional expectation, we immediately get that $(a+1)(2p+r-1)=a$ or $a=(2p+r-1)/(2-2p-r)$. Let us make two observations about the martingale. First, $q$ is not a part of the formula, because in the symmetric case the delays do not affect the ruin probabilities. Second, note that because of $p+r\leq 1$ and $r<1$, we have $02-2p-r>$ 0, so no division by zero. The rest is standard. The OST gives us
$0={\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=A\alpha-B\beta+\frac{2p+r-1}{2-2p-r}(\alpha-\beta).$
Together with $\alpha+\beta=1$ it leads us to
$\alpha=\frac{B-1+\frac{1}{2-2p-r}}{A+B-2+\frac{2}{2-2p-r}}.$ (7)
As a check, let us take a look at some special cases:
  1. 1.
    if $r=0$ (the state 0 is ignored), then we have formula (2),
  2. 2.
    if $12p+r=$ 1 or $p=1-p-r$ (almost the iid case because of an additional symmetry), then $S_{n}$ itself a martingale, and $\alpha=B/(A+B)$
  3. 3.
    if $A=B$ (symmetry), then $\alpha=1/2$.
In fact, because the delays do not really affect the ruin probabilities, the gambler’s ruin problem for a symmetric CRW with delays is equivalent to the gambler’s ruin problem for a two-state symmetric CRW with the probability of staying in a state that is equal to $p+r/2$.
3.2 Expected Waiting Time until Ruin for Symmetric CRW with Delays
Finding ${\mathbf{E}}(\tau)$ is more challenging, because the delays play a significant role. At this point, we have some experience in building “useful” martingales, and this experience suggests the following candidate:
$M_{n}=S_{n}^{2}+aS_{n}X_{n}+bX_{n}^{2}-cn.$
If we are successful in finding the right constants $a$, $b$, and $c$, the OST immediately gives us a formula for the expected waiting time until ruin:
${\mathbf{E}}(\tau)=\frac{1}{c}\big{[}A^{2}\alpha+B^{2}\beta+a(A\alpha+B\beta)+b-{\mathbf{E}}(M_{1})\big{]}.$ (8)
So, the first step is to figure out the conditional expectation ${\mathbf{E}}(X_{n}^{2}|{\mathcal{F}}_{n})$. Using the expressions for indicators, we have that
${\mathbf{E}}(X_{n}^{2}|{\mathcal{F}}_{n-1})=$ $\mathbin{\hphantom{+}}\big{[}1-r\big{]}(X_{n-1}^{2}+X_{n-1})/2$
$+\big{[}1-q\big{]}(1-X_{n-1}^{2})$
$+\big{[}1-r\big{]}(X_{n-1}^{2}+X_{n-1})/2$
$=$ $\mathbin{\hphantom{+}}(1-q)+(q-r)X_{n-1}^{2}.$
Next, we get that
${\mathbf{E}}(M_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}((S_{n-1}+X_{n})^{2}+a(S_{n-1}+X_{n})X_{n}+bX_{n}^{2}-cn|{\mathcal{F}}_{n-1})$
$=$ ${\mathbf{E}}(S_{n-1}^{2}+2S_{n-1}X_{n}+X_{n}^{2}+aS_{n-1}X_{n}+aX_{n}^{2}+bX_{n}^{2}-cn|{\mathcal{F}}_{n-1})$
$=$ $S_{n-1}^{2}+(2+a)(2p+r-1)S_{n-1}X_{n-1}$
$\quad\quad+(1+a+b)(q-r)X_{n-1}^{2}$
$\quad\quad+(1+a+b)(1-q)-cn.$
To make it a martingale $a$, $b$ and $c$ have to satisfy the following linear system:
$\left\{\begin{array}[]{l}(2+a)(2p+r-1)=a\\ (1+a+b)(q-r)=b\\ (1+a+b)(1-q)=c\end{array}\right..$
That leads us to
$a=\frac{2(2p+r-1)}{2-2p-r},\quad\quad b=\frac{(q-r)(2p+r)}{(1+r-q)(2-2p-r)},\quad\quad c=\frac{(1-q)(2p+r)}{(1+r-q)(2-2p-r)}.$
Note that because of the restrictions on $p$, $q$, and $r$, there is no division by zero. The last step is to calculate ${\mathbf{E}}(M_{1})$. Since ${\mathbf{E}}(X_{1}^{2})=(1-q)/(1+r-q)$, we get
${\mathbf{E}}(M_{1})=$ ${\mathbf{E}}(S_{1}^{2}+aS_{1}X_{1}+bX_{1}^{2}-c)$
$=$ ${\mathbf{E}}(X_{1}^{2}+aX_{1}X_{1}+bX_{1}^{2}-c)$
$=$ $(1+a+b)\frac{1-q}{1+r-q}-c$
Finally, formula (8) leads us to
${\mathbf{E}}(\tau)=AB\frac{(1+r-q)(2-2p-r)}{(1-q)(2p+r)}+(A+B)\frac{(2p+r-1)(1+r-q)}{(1-q)(2p+r)}+\frac{(q-r)r}{(1-q)(1+r-q)}.$ (9)
As always, let us consider special cases:
  1. 1.
    if $r=0$ (the state 0 is ignored), then $q$ will be canceled out, and we get formula (4),
  2. 2.
    if $12p+r=$ 1 or $p=1-p-r$ (balanced switches between 1 and -1), then
    ${\mathbf{E}}(\tau)=AB\frac{1+r-q}{1-q}+\frac{(q-r)r}{(1-q)(1+r-q)},$
  3. 3.
    if $12p+r=$ 1 and $r=q$ (the iid symmetric random walk with delays), then ${\mathbf{E}}(\tau)=AB/(1-q)$, a formula from Gut (2013).
3.3 Ruin Probability for General CRW with Delays
Finding closed-form formulas in a general case might be a challenging task. The transition matrix has 6 free parameters. However, it is possible to develop a method that will work, at least numerically, for almost any Markov chain.
In this section $\{X_{n}\}_{n\geq 1}$ is a three-state Markov Chain with state set $\{1,0,-1\}$ and transition matrix
$\begin{pmatrix}p&p_{0}&1-p-p_{0}\\ r_{1}&r&1-r-r_{1}\\ 1-q-q_{0}&q_{0}&q\\ \end{pmatrix},$
where $0\leq p,p_{0},q,q_{0},r,r_{1}\leq 1$ and $0\leq p+p_{0},r+r_{1},q+q_{0}\leq 1$. We will drop here the assumption that the initial distribution is stationary. The initial distribution is given by $(\pi_{1},\pi_{0},1-\pi_{0}-\pi_{1})$, where $0\leq\pi_{1},\pi_{0}\leq 1$ and $0\leq\pi_{1}+\pi_{0}\leq 1$. We will employ the same notation for random walks, ruin probabilities, filtration and stopping time.
The candidate for a martingale in the general case is given by
$M_{n}=\lambda^{S_{n}+aX_{n}^{2}+bX_{n}}.$
If one can find $\lambda\neq 1$, $a$ and $b$ that make $M_{n}$ a martingale, then the OST gives us
${\mathbf{E}}(M_{1})={\mathbf{E}}(M_{\tau})=\lambda^{A+a+b}\alpha+\lambda^{-B+a-b}(1-\alpha).$
Therefore, we have
$\alpha=\frac{{\mathbf{E}}(M_{1})-\lambda^{-B+a-b}}{\lambda^{A+a+b}-\lambda^{-B+a-b}}.$ (10)
So, let us try to find such $\lambda$, $a$ and $b$. First, note that
${\mathbf{E}}(M_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}\left[\lambda^{S_{n-1}+aX_{n-1}^{2}+bX_{n-1}+a(X_{n}^{2}-X_{n-1}^{2})+(b+1)X_{n}-bX_{n-1}}|{\mathcal{F}}_{n-1}\right]$
$=$ $M_{n-1}{\mathbf{E}}\left[\lambda^{a(X_{n}^{2}-X_{n-1}^{2})+(b+1)X_{n}-bX_{n-1}}|{\mathcal{F}}_{n-1}\right].$
Thus we need $\lambda$, $a$ and $b$ that make the conditional expectation of
$Y_{n}=\lambda^{a(X_{n}^{2}-X_{n-1}^{2})+(b+1)X_{n}-bX_{n-1}}$
to be equal to 1. Now, our usual tricks lead us to
${\mathbf{E}}(Y_{n}|{\mathcal{F}}_{n-1})=$ ${\mathbf{E}}(Y_{n}|X_{n-1}=1)\mathbb{I}_{X_{n-1}=1}+{\mathbf{E}}(Y_{n}|X_{n-1}=0)\mathbb{I}_{X_{n-1}=0}+{\mathbf{E}}(Y_{n}|X_{n-1}=-1)\mathbb{I}_{X_{n-1}=-1}$
[…]